OpenStudy (sh3lsh):

Difficult DE Question: Determine all values of x0 and y0 such that the IVP for the the DE given below with initial condtions y(x0) = y0 is guaranteed to have a unique solution. Then, confirm the prediction by finding and analyzing the general solution

OpenStudy (sh3lsh):

\[\frac{ dy }{ dx } = \frac{ 1+\cos(x) }{2-\sin(y) }\]

OpenStudy (anonymous):

The uniqueness theorem should apply, for part one, check if the R.H.S and the derivative of the right hand side w.r.t. y are continuous on a region containing the IC.

OpenStudy (sh3lsh):

How can I automatically tell if this is continuous: http://www.wolframalpha.com/input/?i=d%2Fdy+of+(1%2Bcos(x))%2F(2-sin(y)) It just seems to make things a bit more complicated.

OpenStudy (anonymous):

let \[f(x,y) := \frac{1+\cos(x)}{2-\sin(y)}\] then f(x,y) can only fail to be cont. if ?

OpenStudy (anonymous):

remember, sin and cos are continuous functions.

OpenStudy (sh3lsh):

I don't understand. Am I being daft? It fails to be continuous where this denominator would equate to 0. However, arcsin(2) isn't a thing.

OpenStudy (anonymous):

exactly! but \[|\sin(x)|<1\]

OpenStudy (anonymous):

so it's continuous

OpenStudy (anonymous):

now \[f_y\]

OpenStudy (sh3lsh):

fy? Partial with respect to y?

OpenStudy (sh3lsh):

Yeah, so the fy is always continuous. This means we'll have a general solution wherever. Let's tell for uniqueness by doing a partial w.r.t y. Correct?

OpenStudy (sh3lsh):

Would I have to do this partial? It doesn't seem fun to do by hand. Is there any trick?

OpenStudy (anonymous):

sry, was afk.

OpenStudy (anonymous):

Yes, partial wrt y.

OpenStudy (sh3lsh):

Ugh, well, if somehow with divine intervention and lots of algebra, I obtained this, I would see that this is well continuous because the denominator is never going to be 0. (see Wolfram Alpha link above). I would state there is a unique solution.

OpenStudy (sh3lsh):

Correct?

OpenStudy (anonymous):

yep. Is it really that messy to do?

OpenStudy (sh3lsh):

Actually no, I should have just attempted it.

OpenStudy (sh3lsh):

But how does this general solution of \[2y + \cos(y)=x+\sin(x)\] show anything?

OpenStudy (anonymous):

that's how it is sometimes, and others, it's messy!

OpenStudy (sh3lsh):

Solved simply by separating and integrating with respect to variable.

OpenStudy (anonymous):

I'm not sure, it's exist and is smooth.

OpenStudy (anonymous):

the original analysis gives all the important answers.

OpenStudy (sh3lsh):

You're the best with Differential Equations! Thanks.

OpenStudy (anonymous):

np